# SU(3) symmetry analysis in charmed baryon two body decays with penguin diagram contribution

### v1

@@ I Introduction
In the past decade, the study of baryon decays has attracted increasing attention in particle physics, especially in heavy flavor physics. With the development of experimental facilities like LHC, SuperKEKB, and BEPC, an increasing number of baryon decay channels and new baryon states are being measured for the first time Belle:2013htj ; LHCb:2015yax ; BESIII:2015bjk ; BESIII:2015ysy ; LHCb:2017iph ; LHCb:2017uwr ; Belle:2017ext ; BESIII:2019nep ; LHCb:2024lao ; Belle:2024cmc . A large amount of experimental data has greatly promoted corresponding theoretical research since it provides a new platform to precisely test the Standard Model (SM). Recently, there have been many interesting and influential works on the characters Liu:2022igi ; Tan:2023opd ; Han:2023xbl ; Liu:2023pwr ; Su:2024lzy ; Han:2024ucv ; Peng:2024pyl , decay behaviors Li:2017ndo ; Zhang:2021oja ; Geng:2022osc ; Miao:2022bga ; Aliev:2022tvs ; Rui:2023fiz ; Wang:2024oyi ; Shen:2024aba and angular distributions of baryon Li:2022nim ; Fu:2023ose .
@@
In the analysis of CBTD, SU(3) symmetry serves as a powerful tool that has been extensively utilized in the study of heavy hadrons Lu:2016ogy ; Zhao:2018mov ; Geng:2019xbo ; Jia:2019zxi ; Huang:2021aqu ; Xing:2023dni ; Geng:2023pkr ; Zhong:2024qqs . In the absence of detailed dynamic understanding, the effectiveness of the imposed symmetry can be assessed by how well the known data fit, as indicated by a good $\chi^{2}$ per degree of freedom ($\chi^{2}/d.o.f$). Including the experimental measurement in this year, the number of experimental data is increased to 35 shown in Tables. 5 and 4. This allows us to revisit the global analysis of CBTD, incorporating the contribution from the penguin diagram.
@@ II Theoretical framework including the penguin operator in SU(3) analysis
Based on the above decomposition relation, we can derive the matrix of Hamiltonian in IRA corresponding to each operator in Eq. 2. For the Cabibbo-allowed operator induced by $c\to s\bar{d}u$, the IRA Hamiltonian are
@@
We found that the terms $(H_{15})^{11}_{1}$ and $(H{3})^{1}$ are both proportional to $\lambda_{b}$, introducing a new weak phase. When the operators $c\to ud\bar{d}/s\bar{s}$ contribute to CBTD processes togather, this new phase from $\lambda_{b}$ emerges naturally, particularly in final state interaction scenarios where nearly all Cabibbo-suppressed processes can introduce $\lambda_{b}$ He:2024pxh . Traditionally, terms proportional to $\lambda_{b}$ are neglected due to its small value, $\lambda_{b}\sim O(10^{-4})\ll\lambda_{d,s}$. However, to incorporate the contribution of the penguin operator, terms proportional to $\lambda_{b}$ in $O_{1,2}$ must be included.
@@
One can find that the both $O_{1,2}$ and penguin operator are contribute to the $H_{3}$ and in principle the souces of $H_{3}$ can not be distinguished. Therefore, in our work, we will only estimate the contribution of the $H_{3}$ without distinguish its source.
@@
Here for different ways of contraction, we assign an undetermined parameter, respectively. By absorbing the CKM matrix element $V_{cb}^{*}V_{ub}$ into SU(3) amplitudes $a_{3},b_{3},c_{3},d_{3}$, the $H_{3}$ in the penguin operator can be defined as $(H_{3})^{1}=1$. For the $H_{15}$, since the $\lambda_{b}$ is appeared with $\lambda_{d}-\lambda_{s}$ together in $(H_{15})^{13}_{3},(H_{15})^{31}_{3},(H_{15})^{12}_{2},(H_{15})^{21}_{2}$, one can omit the $\lambda_{b}$ in there matrix element and set $-\lambda_{d}=\lambda_{s}=\sin\theta=0.2265$. Then all the amplitudes of anti-triplet charmed baryon two-body decays can be expressed by these SU(3) amplitudes as shown in Tables. 1 and 2.
@@ III Global analysis of charmed baryon two-body decays with penguin operator
Along with a significant amount of experimental data in Tables 4 and 5, it is imperative to revisit the global analysis of SU(3) irreducible amplitudes and further determine the contribution from the penguin diagrams. Since the experimental measurements for branching ratios of $\Xi_{c}^{0}$ are obtained in terms of the ratios between $Br(\Xi_{c}\to BM)$ and $Br(\Xi_{c}\to\Xi^{-}\pi^{+})$, we use these ratios as data for the fit directly. In Tables 4 and 5, it is evident that among the measured processes, only $\Xi_{c}^{0}\to\Xi^{-}K^{+}$ depends on the amplitude $c_{3}$. However, with only one measurement, $Br(\Xi_{c}^{0}\to\Xi^{-}K^{+})/Br(\Xi_{c}^{0}\to\Xi^{-}\pi^{+})=0.0275(57)$, the corresponding two form factors cannot be determined. Therefore, the amplitude $c_{3}$ cannot be determined by the present data, and we assume $f_{3}^{c}=g_{3}^{c}=0$ in Eq.(14).
@@
Table 3: The global fit results for the real and complex form factors. Case I means the real case in Eq. 13 and Case II means the complex case with the reduced amplitudes in Eq. 18.
@@
vector(f) $f^{a}_{6}=-0.022(23)$ $f^{b}_{6}=0.011(35)$ $f^{c}_{6}=0.0241(65)$ $\delta f^{b}_{6}=-1.088(813)$ $\delta f^{c}_{6}=0.951(729)$
$f^{d}_{6}=-0.018(31)$ $f^{e}_{15}=0.022(32)$ $\delta f^{d}_{6}=-2.593(735)$ $\delta f^{e}_{15}=-0.534(682)$
$f^{a}_{3}=-0.0003(38)$ $f^{b}_{3}=-0.01(2.84)$ $\delta f^{a}_{3}=1.695(811)$ $\delta f^{b}_{3}=-3.042(1.371)$
axial-vector(g) $g^{a}_{6}=0.055(46)$ $g^{b}_{6}=-0.253(77)$ $\delta g^{a}_{6}=1.268(817)$ $\delta g^{b}_{6}=-2.624(0.829)$
$g^{c}_{6}=0.008(43)$ $g^{d}_{6}=-0.044(74)$ $\delta g^{c}_{6}=0.942(655)$ $\delta g^{d}_{6}=-2.368(728)$
$g^{e}_{15}=-0.055(71)$ $\delta g^{e}_{15}=2.501(749)$
$g^{a}_{3}=0.0121(90)$ $g^{b}_{3}=-0.02(18.08)$ $\delta g^{a}_{3}=0.601(828)$ $\delta g^{b}_{3}=2.805(724)$
@@ IV Global analysis of charmed baryon two-body decays with strong phases
Actually, according to the KPW theorem Korner:1970xq ; Pati:1970fg the number of amplitudes in Eq. 13 can be reduced. Considering only the flavor structure, one can find that there are five different ways to contract the $SU(3)_{f}$ indices for $H_{15}$ in Eq.13: $a_{15}$, $b_{15}$, $c_{15}$, $d_{15}$, $e_{15}$. However, taking into account that the color indices of quarks in $H_{15}$ (baryons) must be (anti)symmetric, these five terms can be consolidated into only one term, namely $e_{15}$eng:2023pkr . Similarly, for $H_{3}$, since only $a_{3}$ and $b_{3}$ contribute to the leading term, the aforementioned amplitudes can be further reduced as follows:
@@
Table 4: The experimental data BESIII:2023uvs ; Belle-II:2024jql ; BESIII:2024sfz ; PDG and predicted values for branching ratios, polarization parameters ($\alpha,\beta,\gamma$) and CP violation with the final states $\eta^{(\prime)}$ for two different fits. Case I means the real case in Eq. 13 and Case II means the complex case with the reduced amplitudes in Eq. 18. Here the data with * mean the ratio of decay width between the corresponding processes and $\Gamma(\Xi_{c}^{0}\to\Xi^{-}\pi^{+})$.
@@
The 14 complex form factors corresponding to the 7 amplitudes can be expressed as
@@
Table 5: The experimental data and predicted values for branching ratios, polarization parameters ($\alpha,\beta,\gamma$) and CP violation for two different fits. Case I means the real case in Eq. 13 and Case II means the complex case with the reduced amplitudes in Eq. 18. Here the data with * mean the ratio of decay width between the corresponding processes and $\Gamma(\Xi_{c}^{0}\to\Xi^{-}\pi^{+})$.

### v3

@@ I Introduction
In the past decade, the study of baryon decays has attracted increasing attention in particle physics, especially in heavy flavor physics. With the development of experimental facilities like LHC, SuperKEKB, and BEPC, an increasing number of baryon decay channels and new baryon states are being measured for the first time Belle:2013htj ; LHCb:2015yax ; BESIII:2015bjk ; BESIII:2015ysy ; LHCb:2017iph ; LHCb:2017uwr ; Belle:2017ext ; BESIII:2019nep ; LHCb:2024lao ; Belle:2024cmc . A large amount of experimental data has greatly promoted corresponding theoretical research since it provides a new platform to precisely test the Standard Model (SM). Recently, there have been many interesting and influential works on the characters Liu:2022igi ; Deng:2023csv ; Su:2024lzy ; Han:2023xbl ; Liu:2023pwr ; Deng:2023qaf ; Tan:2023opd ; Han:2024ucv ; Peng:2024pyl , decay behaviors Li:2017ndo ; Zhang:2021oja ; Geng:2022osc ; Miao:2022bga ; Aliev:2022tvs ; Rui:2023fiz ; Wang:2024oyi ; Shen:2024aba and angular distributions of baryon Li:2022nim ; Fu:2023ose .
@@
In the analysis of CBTD, SU(3) symmetry serves as a powerful tool that has been extensively utilized in the study of heavy hadrons Lu:2016ogy ; Zhao:2018mov ; Geng:2019xbo ; Jia:2019zxi ; Wang:2020gmn ; Huang:2021aqu ; Xing:2023dni ; Geng:2023pkr ; Zhong:2024qqs ; Wang:2024ztg . In the absence of detailed dynamic understanding, the effectiveness of the imposed symmetry can be assessed by how well the known data fit, as indicated by a good $\chi^{2}$ per degree of freedom ($\chi^{2}/d.o.f$). Including the experimental measurement in this year, the number of experimental data is increased to 35 shown in Tables. 5 and 4. This allows us to revisit the global analysis of CBTD, incorporating the contribution from the penguin diagram.
@@ II Theoretical framework including the penguin operator in SU(3) analysis
It is noted that in previous work, the decomposition has been presented in an alternative form He:2000ys ; Wang:2020gmn . We demonstrate in Appendix A that the two decomposition forms are equivalent.
Based on the aforementioned decomposition relationship, we can derive the Hamiltonian matrix in IRA corresponding to each operator in Eq. 2. For the Cabibbo-allowed operator induced by $c\to s\bar{d}u$, the IRA Hamiltonian are
@@
We found that the terms $(H_{15})^{11}_{1}$ and $(H_{3})^{1}$ are both proportional to $\lambda_{b}$, introducing a new weak phase. When the operators $c\to ud\bar{d}/s\bar{s}$ contribute to CBTD processes together, this new phase from $\lambda_{b}$ emerges naturally, particularly in final state interaction scenarios where nearly all Cabibbo-suppressed processes can introduce $\lambda_{b}$ He:2024pxh . Traditionally, terms proportional to $\lambda_{b}$ are neglected due to its small value, $\lambda_{b}\sim O(10^{-4})\ll\lambda_{d,s}$. However, to incorporate the contribution of the penguin operator, terms proportional to $\lambda_{b}$ in $O_{1,2}$ must be included.
@@
One can find that the both $O_{1,2}$ and penguin operator contribute to the $H_{3}$ and in principle the souces of $H_{3}$ can not be distinguished. Therefore, in our work, we will only estimate the contribution of the $H_{3}$ without distinguish its source.
@@
Here for different ways of contraction, we assign an undetermined parameter, respectively. By absorbing the CKM matrix element $V_{cb}^{*}V_{ub}$ into SU(3) amplitudes $a_{3},b_{3},c_{3},d_{3}$, the $H_{3}$ in the penguin operator can be defined as $(H_{3})^{1}=1$. For the $H_{15}$, since the $\lambda_{b}$ appears with $\lambda_{d}-\lambda_{s}$ together in $(H_{15})^{13}_{3},(H_{15})^{31}_{3},(H_{15})^{12}_{2},(H_{15})^{21}_{2}$, one can omit the $\lambda_{b}$ in these matrix elements and set $-\lambda_{d}=\lambda_{s}=\sin\theta=0.2265$. Then all the amplitudes of anti-triplet charmed baryon two-body decays can be expressed by these SU(3) amplitudes as shown in Tables. 1 and 2.
@@ III Global analysis of charmed baryon two-body decays with penguin operator
Along with a significant amount of experimental data in Tables 4 and 5, it is imperative to revisit the global analysis of SU(3) irreducible amplitudes and further determine the contribution from the penguin diagrams. Since the experimental measurements for branching ratios of $\Xi_{c}^{0}$ are obtained in terms of the ratios between $Br(\Xi_{c}\to BM)$ and $Br(\Xi_{c}\to\Xi^{-}\pi^{+})$, we use these ratios as data for the fit directly. In Tables 4 and 5, it is evident that among the measured processes, only $\Xi_{c}^{0}\to\Xi^{-}K^{+}$ depends on the amplitude $c_{3}$. However, with only one measurement, $Br(\Xi_{c}^{0}\to\Xi^{-}K^{+})/Br(\Xi_{c}^{0}\to\Xi^{-}\pi^{+})=0.0275(57)$, the corresponding two form factors cannot be determined. Therefore, the amplitude $c_{3}$ cannot be determined by the present data. Fortunately, the KPW theorem Korner:1970xq ; Pati:1970fg requires that the color structure of the two-quark operators in the effective Hamiltonians is symmetric, which conflicts with the antisymmetric color structure of quarks in the baryon state. Since, in $c_{3}$ term in Eq. (13), the two-quark fields in the charmed baryon are directly connected to the Hamiltonian $H_{3}^{i}$, the contribution of $c_{3}$ is suppressed by the color symmetry. Therefore, in our work, we assume $f_{3}^{c}=g_{3}^{c}=0$ in Eq. (14).
@@
Our predictions reveal that the decay modes involving $\eta$ and $\eta^{\prime}$ in the final state exhibit different behaviors. Taking the $\Lambda_{c}^{+}$ decays as an example, the branching ratios of the $\Lambda_{c}^{+}\to\Sigma^{+}$ modes in both cases are comparable, whereas those of the $\Lambda_{c}^{+}\to p$ modes differ significantly. This intriguing phenomenon also manifests in $\Xi_{c}$ decay modes. These distinct behaviors can be attributed to the fitted form factors in Table. 3, which show that the largest values are $g_{a}=0.126(10)$, $g_{a\prime}=0.160(32)$ and $g_{6}^{b}=-0.186(14)$. Since $\eta$ and $\eta^{\prime}$ are mixed states of $\eta_{q}$ and $\eta_{s}$, the amplitudes of charmed baryon decay modes involving $\eta$ and $\eta^{\prime}$ in the final state can be expressed as a combination of the decay amplitudes with $\eta_{q}$ and $\eta_{s}$, parameterized by the mixing angle $\phi=(39.3\pm 1.0)^{\circ}$, as shown in Table. 2. In the case of the $\Lambda_{c}^{+}\to\Sigma^{+}\eta_{q/s}$ modes, the cancellation between the form factors $g^{a}=g_{6}^{a}-g_{15}^{a}$ and $g^{6}_{b}$ reduces the $\Lambda_{c}^{+}\to\Sigma^{+}\eta_{q}$ decay amplitude, making the $\Lambda_{c}^{+}\to\Sigma^{+}\eta_{s}$ amplitude the dominant contribution. Consequently, this leads to comparable branching ratios for the $\Lambda_{c}^{+}\to\Sigma^{+}$ modes. In the $\Lambda_{c}^{+}\to p\eta_{q/s}$ modes, the dominant contributions come from the amplitudes $a_{3}$, $b_{3}$, and $d_{3}$, making the $\Lambda_{c}^{+}\to p\eta_{q/s}$ amplitudes comparable. Consequently, the two modes with $\eta_{q/s}$ interfere constructively and destructively in the amplitude of $\Lambda_{c}^{+}\to p\eta^{(\prime)}$, leading to significantly different branching ratios.
For the polarization parameters $\alpha$, the situation becomes significantly more complex because they reflect the discrepancies between P-conservation and P-violation factors. Consequently, the predictions we offer tend to have greater uncertainties. Thus, it is most appropriate to address this topic once sufficient experimental data is available, enabling more accurate results that account for complex phase factors.
@@
Table 3: The global fit results for the real (Case I in Eq. 13) and complex form factors (Case II in Eq. 19).
@@
vector(f) $f^{a}_{6}=0.022(23)$ $f^{b}_{6}=0.023(51)$ $f^{c}_{6}=0.0415(87)$ $\delta f^{b}_{6}=-1.088(813)$ $\delta f^{c}_{6}=0.951(729)$
$f^{d}_{6}=-0.018(31)$ $f^{e}_{15}=0.026(41)$ $\delta f^{d}_{6}=-2.593(735)$ $\delta f^{e}_{15}=-0.534(682)$
$f^{a}_{3}=0.0024(55)$ $f^{b}_{3}=0.01(14.29)$ $\delta f^{a}_{3}=1.695(811)$ $\delta f^{b}_{3}=-3.042(1.371)$
axial-vector(g) $g^{a}_{6}=0.184(67)$ $g^{b}_{6}=0.29(11)$ $\delta g^{a}_{6}=1.268(817)$ $\delta g^{b}_{6}=-2.624(0.829)$
$g^{c}_{6}=0.014(54)$ $g^{d}_{6}=0.062(99)$ $\delta g^{c}_{6}=0.942(655)$ $\delta g^{d}_{6}=-2.368(728)$
$g^{e}_{15}=0.069(97)$ $\delta g^{e}_{15}=2.501(749)$
$g^{a}_{3}=0.015(13)$ $g^{b}_{3}=0.02(24.13)$ $\delta g^{a}_{3}=0.601(828)$ $\delta g^{b}_{3}=2.805(724)$
@@ IV Global analysis of charmed baryon two-body decays with strong phases
Actually, according to the KPW theorem Korner:1970xq ; Pati:1970fg the number of amplitudes in Eq. 13 can be reduced. To illustrate the flavor and color symmetries, the Hamiltonian can be decomposed into symmetric and antisymmetric components, using the current-current operator as an example, He:2024pxh
$\mathcal{H}_{eff}=\frac{G_{F}}{\sqrt{2}}\sum_{\lambda=\pm}C_{\lambda}(\mathcal{H}_{\lambda})^{ij}_{k}[\bar{q}_{i}q^{k}]_{V-A}[\bar{q}_{j}c]_{V-A},$ (18)
where $C_{\pm}=(C_{1}\pm C_{2})/2$ and $(H_{+/-})^{ij}_{k}$ is the symmetry/anti-symmetry for the indices $i,j$. The symmetry or antisymmetry of the anti-quark field also reflects the symmetry of its color. Considering only the flavor structure, one can find that there are five different ways to contract the $SU(3)_{f}$ indices for $H_{15}$ in Eq.13: $a_{15}$, $b_{15}$, $c_{15}$, $d_{15}$, $e_{15}$. However, taking into account that color indices of quarks in $H_{15}$ (baryons) must be (anti)symmetric, these five terms can be consolidated into only one term, namely $e_{15}$ Geng:2023pkr . Similarly, for $H_{3}$, since only $a_{3}$ and $b_{3}$ contribute to the leading term, the aforementioned amplitudes can be further reduced as follows:
@@
We also observe that the anti-symmetric part of the Hamiltonian in Eq. (18) violates flavor symmetry, particularly in the $c\to d\bar{s}u$ processes. This also leads to the breaking of isospin symmetry. Using the isospin-symmetric processes $(\Xi_{c}^{0}\to p\pi^{-},\Xi_{c}^{+}\to n\pi^{+})$ and $(\Xi_{c}^{0}\to n\pi^{0},\Xi_{c}^{+}\to p\pi^{0})$ as examples, the SU(3) amplitude $c_{6}$, corresponding to the anti-symmetric Hamiltonian, along with $c_{15}$ and $d_{15}$, corresponding to the symmetric part of the Hamiltonian, together break the isospin symmetry, as shown in Table. 1. As a result of this isospin symmetry violation, the predicted branching ratios and polarization parameters in Case I of Table. 4 differ significantly. After omitting the amplitudes $c_{15}$ and $d_{15}$ (Case II), one can derive the isospin symmetry relation as follows:
$\mathcal{M}(\Xi_{c}^{0}\to p\pi^{-})=\mathcal{M}(\Xi_{c}^{+}\to n\pi^{+})$
$=-\sqrt{2}\mathcal{M}(\Xi_{c}^{0}\to n\pi^{0})=\sqrt{2}\mathcal{M}(\Xi_{c}^{+}\to p\pi^{0}).$ (20)
The situation also applies to the isospin-symmetric processes $(\Xi_{c}^{0}\to n\eta,\;\Xi_{c}^{+}\to p\eta)$ and $(\Xi_{c}^{0}\to n\eta^{\prime},\Xi_{c}^{+}\to p\eta^{\prime})$. One can similarly derive their isospin symmetry relations after omitting the SU(3) amplitudes corresponding to the symmetric part of the Hamiltonian, as follows:
$\mathcal{M}(\Xi_{c}^{0}\to n\eta)=\mathcal{M}(\Xi_{c}^{+}\to p\eta),$
$\mathcal{M}(\Xi_{c}^{0}\to n\eta^{\prime})=\mathcal{M}(\Xi_{c}^{+}\to p\eta^{\prime}).$ (21)
Therefore, in case II, the KPW theorem necessitates neglecting the color-suppressed SU(3) amplitudes $a_{15},b_{15},c_{15},d_{15}$. This ensures that the isospin symmetry of the aforementioned decay channels will be naturally satisfied.
Table 4: The experimental data BESIII:2023uvs ; Belle-II:2024jql ; BESIII:2024sfz ; PDG and predicted values for branching ratios, polarization parameters ($\alpha,\beta,\gamma$) and CP violation with the final states $\eta^{(\prime)}$ for two different fits: real case (Case I in Eq. 13) and complex case ( Case II means in Eq. 19). Here the data with * mean the ratio of decay width between the corresponding processes and $\Gamma(\Xi_{c}^{0}\to\Xi^{-}\pi^{+})$.
@@
Considering the 7 SU(3) amplitude, the 14 complex form factors can be expressed as
@@
Table 5: The experimental data and predicted values for branching ratios, polarization parameters ($\alpha,\beta,\gamma$) and CP violation for two different fits. The details are the same with Table. 4.
@@ Acknowledgements
Appendix A The decomposition of IRA heavy-to-light Hamiltonian
There is another form of decomposition of the IRA heavy-to-light Hamiltonian as
$H^{ij}_{k}$ $=$ $\frac{1}{8}(H_{15})^{ij}_{k}+\frac{1}{4}\epsilon^{ijl}(H_{\bar{6}})_{lk}+\delta^{j}_{k}\left(\frac{3}{8}(H^{t}_{3})^{i}-\frac{1}{8}(H^{p}_{3})^{i}\right)$ (25)
$+\delta^{i}_{k}\left(\frac{3}{8}(H^{p}_{3})^{j}-\frac{1}{8}(H^{t}_{3})^{j}\right),$
where $(H^{t}_{3})^{i}=H^{im}_{m}$ and $(H^{p}_{3})^{i}=H^{mi}_{m}$. At first glance, we might naively say that the expression of the decomposition is very different from our formula in Eq. 4. Actually, the two expressions are equivalent, as can be clearly understood from the operator aspect of the IRA Hamiltonian. In this aspect, the Hamiltonian matrix $H^{ij}_{k}$ is expressed as
$O(H)^{ij}_{k}=[\bar{q}_{j}c][\bar{q}_{i}q_{k}],\quad(q_{1},q_{2},q_{3})=(u,d,s),$ (26)
where we omit the color and Lorentz structure in SU(3) symmetry. Correspondingly, the matrix elements of $H_{15}$, $H_{\bar{6}}$, and $H_{3}$ both can be obtained. By using the transformation relations of $H_{3/{3\prime}}$ and $H_{3}^{t/p}$
$(H_{3})^{ij}_{k}$ $=$ $\delta^{i}_{k}(H^{t}_{3})^{j}+\delta^{j}_{k}(H^{p}_{3})^{i},$
$(H_{3\prime})^{ij}_{k}$ $=$ $\delta^{i}_{k}(H^{p}_{3})^{j}+\delta^{j}_{k}(H^{t}_{3})^{i},$ (27)
one can directly find that the $3$ representations in Eq. 4 and Eq. 25 are equivalent. And the the matrix elements $H_{3}^{t/p}$ are
$O(H_{3}^{p})^{1}$ $=$ $[\bar{u}c][\bar{u}u]+[\bar{u}c][\bar{d}d]+[\bar{u}c][\bar{s}s],$
$O(H_{3}^{p})^{2}$ $=$ $[\bar{d}c][\bar{u}u]+[\bar{d}c][\bar{d}d]+[\bar{d}c][\bar{s}s],$
$O(H_{3}^{p})^{3}$ $=$ $[\bar{s}c][\bar{u}u]+[\bar{s}c][\bar{d}d]+[\bar{s}c][\bar{s}s],$
$O(H_{3}^{t})^{1}$ $=$ $[\bar{u}c][\bar{u}u]+[\bar{d}c][\bar{u}d]+[\bar{s}c][\bar{u}s],$
$O(H_{3}^{t})^{2}$ $=$ $[\bar{u}c][\bar{d}u]+[\bar{d}c][\bar{d}d]+[\bar{s}c][\bar{d}s],$
$O(H_{3}^{t})^{3}$ $=$ $[\bar{u}c][\bar{s}u]+[\bar{d}c][\bar{s}d]+[\bar{s}c][\bar{s}s].$ (28)
Besides, the matrix elements of $H_{\bar{6}}$ and $H_{15}$ are
$O(H_{\bar{6}})^{11}$ $=$ $[\bar{d}c][\bar{s}u]-[\bar{s}c][\bar{d}u],\;\;O(H_{\bar{6}})^{22}=[\bar{s}c][\bar{u}d]-[\bar{u}c][\bar{s}d],\;\;O(H_{\bar{6}})^{33}=[\bar{u}c][\bar{d}s]-[\bar{d}c][\bar{u}s],$
$O(H_{\bar{6}})^{12}$ $=$ $\frac{1}{2}([\bar{s}c][\bar{u}u]-[\bar{u}c][\bar{s}u]+[\bar{d}c][\bar{s}d]-[\bar{s}c][\bar{d}d]),\;\;O(H_{\bar{6}})^{23}=\frac{1}{2}([\bar{u}c][\bar{d}d]-[\bar{d}c][\bar{u}d]+[\bar{s}c][\bar{u}s]-[\bar{u}c][\bar{s}s]),$
$O(H_{\bar{6}})^{31}$ $=$ $\frac{1}{2}([\bar{d}c][\bar{s}s]-[\bar{s}c][\bar{d}s]+[\bar{u}c][\bar{d}u]-[\bar{d}c][\bar{u}u]),$
$O(H_{15})^{11}_{1}$ $=$ $[\bar{u}c][\bar{u}u]-\frac{1}{2}([\bar{u}c][\bar{d}d]+[\bar{d}c][\bar{u}d]+[\bar{s}c][\bar{u}s]+[\bar{u}c][\bar{s}s]),\quad O(H_{15})^{23}_{1}=[\bar{s}c][\bar{d}u]+[\bar{d}c][\bar{s}u]$
$O(H_{15})^{22}_{2}$ $=$ $[\bar{d}c][\bar{d}d]-\frac{1}{2}([\bar{u}c][\bar{d}u]+[\bar{s}c][\bar{d}s]+[\bar{d}c][\bar{u}u]+[\bar{d}c][\bar{s}s]),\quad O(H_{15})^{13}_{2}=[\bar{u}c][\bar{s}d]+[\bar{s}c][\bar{u}d]$
$O(H_{15})^{33}_{3}$ $=$ $[\bar{s}c][\bar{s}s]-\frac{1}{2}([\bar{u}c][\bar{s}u]+[\bar{d}c][\bar{d}s]+[\bar{s}c][\bar{u}u]+[\bar{s}c][\bar{d}d]),\quad O(H_{15})^{12}_{3}=[\bar{d}c][\bar{u}s]+[\bar{u}c][\bar{d}s]$
$O(H_{15})^{21}_{1}$ $=$ $\frac{3}{4}([\bar{d}c][\bar{u}u]+[\bar{u}c][\bar{d}u])-\frac{1}{2}[\bar{d}c][\bar{d}d]-\frac{1}{4}([\bar{s}c][\bar{d}s]+[\bar{d}c][\bar{s}s]),\;O(H_{15})^{11}_{2}=2[\bar{u}c][\bar{u}d],$
$O(H_{15})^{21}_{2}$ $=$ $\frac{3}{4}([\bar{d}c][\bar{u}d]+[\bar{u}c][\bar{d}d])-\frac{1}{2}[\bar{u}c][\bar{u}u]-\frac{1}{4}([\bar{s}c][\bar{u}s]+[\bar{u}c][\bar{s}s]),\;O(H_{15})^{11}_{3}=2[\bar{u}c][\bar{u}s],$
$O(H_{15})^{31}_{1}$ $=$ $\frac{3}{4}([\bar{s}c][\bar{u}u]+[\bar{u}c][\bar{s}u])-\frac{1}{2}[\bar{s}c][\bar{s}s]-\frac{1}{4}([\bar{d}c][\bar{s}d]+[\bar{s}c][\bar{d}d]),\;O(H_{15})^{22}_{3}=2[\bar{d}c][\bar{d}s],$
$O(H_{15})^{31}_{3}$ $=$ $\frac{3}{4}([\bar{s}c][\bar{u}s]+[\bar{u}c][\bar{s}s])-\frac{1}{2}[\bar{u}c][\bar{u}u]-\frac{1}{4}([\bar{d}c][\bar{u}d]+[\bar{u}c][\bar{d}d]),\;O(H_{15})^{22}_{1}=2[\bar{d}c][\bar{d}u],$
$O(H_{15})^{32}_{2}$ $=$ $\frac{3}{4}([\bar{s}c][\bar{d}d]+[\bar{d}c][\bar{s}d])-\frac{1}{2}[\bar{s}c][\bar{s}s]-\frac{1}{4}([\bar{u}c][\bar{s}u]+[\bar{s}c][\bar{u}u]),\;O(H_{15})^{33}_{1}=2[\bar{s}c][\bar{s}u],$
$O(H_{15})^{32}_{3}$ $=$ $\frac{3}{4}([\bar{s}c][\bar{d}s]+[\bar{d}c][\bar{s}s])-\frac{1}{2}[\bar{d}c][\bar{d}d]-\frac{1}{4}([\bar{u}c][\bar{d}u]+[\bar{d}c][\bar{u}u]),\;O(H_{15})^{33}_{2}=2[\bar{s}c][\bar{s}d].$ (29)
Comparing with the previous work He:2000ys ; Wang:2020gmn , one can find that the expressions of $H_{\bar{6}}$ and $H_{15}$ in Eq. 4 and Eq. 25 can be transformed by the factors $-\frac{1}{2}$ and $\frac{1}{4}$, respectively. These two factors exactly compensate for the difference in coefficients of Hamiltonian. Therefore, after considering the differences from both Hamiltonian and the operators, the decomposition formulas in Eq. 4 and Eq. 25 are equivalent to each other.